On the Stabilizability of Systems with Control Delay Via Output Feedback

Jiang Wei University of Anhui, Hefei, China

In this paper, we will discusses the stabilizability of the systems with control delay. Firstly we give a transformation, by which we transform the discussed systems with control delay to be the systems without control delay. We give the relationship of stabilizability of these two kind systems. Then we discuss the stabilizability via output feedback of the the systems without control delay, we give the conditions for the stabilizability of systems with control delay.